Two groups each of three children respectively 2 boys and 1 girl and 1 boy and 2 girls. One child is drawn at random from each group. Calculate the probability that:1.both will be boys
2.1 boy and other girls
1
Expert's answer
2019-04-23T03:42:12-0400

1.We choose a boy from each group independently. Then

P{both will be boys}=P{a child from the first group will be a boy}P{a child from the second group will be a boy}=2/3*1/3=2/9..

2.P{1 boy and other girls}=P{a boy from the first group, a girl from the second group}+

P{a boy from the second group, a girl from the first group}=2/3*2/3+1/3*1/3=5/9.